# Changsha, China

### v1

@@ I Introduction
To bridge this gap, we perform the first systematic study on applying state-of-the-art model editing approaches to repair the inaccuracy of LLMs4Code. To that end, we first build an evaluation benchmark, Code Large Language Models Editing Evaluation (CLMEEval), which consists of two datasets, CoNaLa-Edit (CNLE) and CodeSearchNet-Edit (CSNE), corresponding to the editing of code knowledge in the context of two widely-studied software engineering tasks: a natural language to programming language (NL2PL) code generation task and a programming language to natural language (PL2NL) code summarization task. Drawing from this benchmark, we employ model editing techniques to rectify the inaccuracies produced by LLMs4Code, mirroring real-world scenarios where updates to code knowledge within a model are necessary. This could involve situations like changes in required APIs for completing specific coding tasks or shifts in the primary functionality of a method due to code changes. Following the common practice in the model editing domain [25], our study evaluates the approaches from four dimensions: Effectiveness, the success rate on editing instances; Generalization, the success rate on tests that are semantically identical to the editing instances; Specificity, the success rate on tests unrelated to the editing instances; and Fluency, the fluency of the contents generated by the model [25].
We select six state-of-the-art model editing approaches from the three categories mentioned above and three widely-used LLMs4Code, i.e., CodeLlama (7B)[26], CodeQwen-1.5 (7B) [27], and StableCode (3B)[28] as our study subjects. Through an extensive evaluation, our study makes the following important findings:
@@
Moreover, through a case analysis on the best-performing editing approach in terms of the effectiveness and specificity, i.e., GRACE [29], we identify its weakness in the inability to distinguish between the semantics of different inputs. Based on this observation, we propose an augmented strategy (named as A-GRACE) where we introduce an encoder to GRACE that allows it to better capture the input semantics via contrastive learning, and thus improves the generalization of GRACE. Results show that A-GRACE significantly improves the generalization while achieves similar effectiveness and specificity compared to GRACE. For instance, on CNLE, A-GRACE improves the generalization of GRACE from almost zero to an average of 80.86% in terms of the Exact Match metric.
@@ IV-A RQ1: Proficiency of Editing Techniques on LLMs4Code
    None of the existing model editing techniques can achieve satisfactory results simultaneously across the effectiveness, generalization, and specificity. GRACE achieves the best overall performance among all the techniques with the highest effectiveness and specificity values. Specifically, its ROUGE-L value can reach nearly one hundred pencentage (i.e., 99.63%) when editing CodeLlama on the PL2NL task. Nonetheless, it shows nearly no improvement in generalization compared to the original model. FT-L achieves good effectiveness and generalization in editing CodeQwen on the CSNE dataset, but its significant decline in specificity and fluency compared to the original model indicates a substantial impact on the non-targeted knowledge and generative capability of the model. ROME exhibits the best performance in editing CodeLlama on the CSNE dataset, but similarly, it significantly impacts the original knowledge of the model. The performances of MEMIT, PMET, and MALMEN are generally poor. On the one hand, they do not effectively incorporate the required knowledge into LLMs4Code, which is revealed by the comparatively low effectiveness. On the other hand, they damage the original knowledge of the model significantly according to the low specificity. In terms of fluency, FT-L produces the overall best results, with sometimes even increasing the fluency of the edited model (e.g., when editing CodeLlama on the CSNE dataset).
@@
    Most model editing techniques perform poorly on LLMs4Code, being far less proficient compared with editing general LLMs[21]. For instance, FT-L achieved 100% effectiveness in terms of the EM when editing Llama-2 on factual knowledge datasets[21], while its effectiveness on both CNLE and CSNE datasets are generally poor, often being less than 65%. ROME, MEMIT, PMET, and MALMEN achieved nearly 100% effectiveness and over 85% generalization without significant decreases in specificity and fluency when editing GPT-J[56] on factual knowledge datasets[24]. However, their effectiveness and generalization on the CNLE and CSNE datasets did not surpass 80%. Additionally, there was a notable decrease in both specificity and fluency when utilizing these model editing techniques. Such results indicate that compared with factual knowledge, editing code-related knowledge is a more challenging task that deserves further investigations in the future.
@@
    Generalization is a universal challenge for state-of-the-art model editing techniques. Except for four instances where FT-L and ROME edit LLMs4Codeusing the CSNE dataset, achieving a generalization rate surpassing 50%, the generalization rate in all other scenarios dips below 50%.
@@ IV-C RQ3: Improvements of Editing Techniques on LLMs4Code
Specifically, we propose A-GRACE, which introduces an MLP encoder composed of two linear layers. Different from GRACE, the encoder utilizes the mean of all token representations as the key $q\in\mathbb{R}^{d_{n}}$, aiming at capturing the information of the entire sequence from a holistic perspective. The encoding process of the MLP encoder for the key can be described as $q^{*}=W_{2}\sigma(W_{1}q)$, where $W_{1}\in\mathbb{R}^{d_{n}\times\frac{d_{n}}{4}}$, $W_{2}\in\mathbb{R}^{\frac{d_{n}}{4}\times d_{o}}$, $d_{o}$ is the output dimension, and $\sigma$ is the activation function. We set it as leaky ReLU [61] to prevent neuron deactivation. After that, we train the MLP encoder using contrastive learning on the training set. The training target is to differentiate between GRI and SRI, ensuring that the distance between the keys of the OI and the GRI is as close as possible, while the distance between the keys of the OI and the SRI is as far as possible. To that end, we use the hidden states of the $<$OI, GRI$>$ pairs as positive samples and the hidden states of the $<$OI, SRI$>$ pairs as negative samples for training, with the following loss function [62]:
@@
w/o mean 80.08 91.61 95.10 65.33 81.79 88.15 1.17 18.13 30.21 473.25 87.90 95.90 98.87 86.30 95.81 98.91 43.73 84.69 87.70 504.32
We set the output dimension of the MLP encoder to $d_{o}=256$ to balance encoding performance and efficiency; the slope of the leaky ReLU to 0.5; the dropout rate to 0.5; the margin to 1 to follow the original GRACE approach. During the training phase, we use a batch size of 64, a learning rate of 1e-4, and a weight decay of 1e-4 for 100 epochs. We apply an early stopping strategy with a patience of 5 and use the checkpoint with the best performance on the test set as the final MLP encoder weights.
Results. We set the other parameters of A-GRACE to be consistent with those of GRACE. The evaluation results on the CLMEEval benchmark are shown in Table VI. The results show that A-GRACE consistently outperforms GRACE on all the metrics, particularly in terms of the generalization. For instance, when editing CodeLlama on the CNLE dataset, the EM in terms of the generalization is increased from 0.05% of the vanilla GRACE to 75.50% of A-GRACE. On the CSNE dataset, this value is boosted from 8.68% to 96.63%.
@@
Additionally, in this table, “w/o CL” refers to the ablation result of directly using the mean of all token representations as the key without using contrastive learning, and “w/o mean” refers to the ablation result of using the MLP encoder trained by contrastive learning to encode the last token representation. The ablation results indicate two major points: first, directly using the mean of all token representations yields almost no improvement in generalization; second, although encoding the last token representation with the MLP encoder trained by contrastive learning can significantly improve the generalization, it also leads to a severe decline in specificity and fluency. This demonstrates the rationale of our approach design, that is, using a contrastive-learning-augmented MLP encoder to encode the mean of all token representations as the key is the optimal choice.
@@
(1) Using the last token representation as the key cannot help GRACE effectively distinguish between GRI and SRI; (2) By introducing an contrastively-trained encoder in GRACE and encoding the mean of all token representations (the proposed A-GRACE), we successfully enhances the generalization of GRACE while neglectably affecting other aspects.

### v2

@@ I Introduction
To bridge this gap, we perform the first systematic study on applying state-of-the-art model editing approaches to repair the inaccuracy of LLMs4Code. To that end, we first build an evaluation benchmark, Code Large Language Models Editing Evaluation (CLMEEval), which consists of two datasets, CoNaLa-Edit (CNLE) and CodeSearchNet-Edit (CSNE), corresponding to the editing of code knowledge in the context of two widely-studied software engineering tasks: a natural language to programming language (NL2PL) code generation task and a programming language to natural language (PL2NL) code summarization task. Drawing from this benchmark, we employ model editing techniques to rectify the inaccuracies produced by LLMs4Code, mirroring real-world scenarios where updates to code knowledge within a model are necessary. This could involve situations like changes in required APIs for completing specific coding tasks or shifts in the primary functionality of a method due to code changes. Following the common practice in the model editing domain [25], our study evaluates the approaches from four dimensions: Effectiveness, the success rate on editing instances; Generalization, the success rate on tests that are semantically identical to the editing instances; Specificity, the success rate on tests unrelated to the editing instances; and Fluency, the fluency of the contents generated by the model [25]. We select six state-of-the-art model editing approaches from the three categories mentioned above and three widely-used LLMs4Code, i.e., CodeLlama (7B)[26], CodeQwen-1.5 (7B) [27], and StableCode (3B)[28] as our study subjects. Through an extensive evaluation, our study makes the following important findings:
@@
Moreover, through a case analysis on the best-performing editing approach in terms of the effectiveness and specificity, i.e., GRACE [29], we identify its weakness in the inability to distinguish between the semantics of different inputs. Based on this observation, we propose an augmented strategy (named as A-GRACE) where we introduce an encoder to GRACE that allows it to better capture the input semantics via contrastive learning, and thus improves the generalization of GRACE. Results show that A-GRACE significantly improves the generalization while achieves similar effectiveness and specificity compared to GRACE. For instance, on CNLE, A-GRACE improves the generalization of GRACE from almost zero to an average of 76.63% in terms of the Exact Match metric.
@@ IV-A RQ1: Proficiency of Editing Techniques on LLMs4Code
    None of the existing model editing techniques can achieve satisfactory results simultaneously across the effectiveness, generalization, and specificity. GRACE achieves the best overall performance among all the techniques with the highest effectiveness and specificity values. Specifically, its ROUGE-L value can reach nearly one hundred pencentage (i.e., 99.29%) when editing CodeLlama on the NL2PL task. Nonetheless, it shows nearly no improvement in generalization compared to the original model. FT-L achieves good effectiveness and generalization in editing CodeQwen on the CSNE dataset, but its significant decline in specificity and fluency compared to the original model indicates a substantial impact on the non-targeted knowledge and generative capability of the model. ROME exhibits the best performance in editing CodeLlama on the CSNE dataset, but similarly, it significantly impacts the original knowledge of the model. The performances of MEMIT, PMET, and MALMEN are generally poor. On the one hand, they do not effectively incorporate the required knowledge into LLMs4Code, which is revealed by the comparatively low effectiveness. On the other hand, they damage the original knowledge of the model significantly according to the low specificity. In terms of fluency, FT-L produces the overall best results, with sometimes even increasing the fluency of the edited model (e.g., when editing CodeLlama on the CSNE dataset).
@@
    Most model editing techniques perform poorly on LLMs4Code, being far less proficient compared with editing general LLMs[21]. For instance, FT-L achieved 100% effectiveness in terms of the EM when editing Llama-2 on factual knowledge datasets[21], while its effectiveness on both CNLE and CSNE datasets are generally poor, often being less than 70%. ROME, MEMIT, PMET, and MALMEN achieved nearly 100% effectiveness and over 85% generalization without significant decreases in specificity and fluency when editing GPT-J[56] on factual knowledge datasets[24]. However, their effectiveness and generalization on the CNLE and CSNE datasets did not surpass 80%. Additionally, there was a notable decrease in both specificity and fluency when utilizing these model editing techniques. Such results indicate that compared with factual knowledge, editing code-related knowledge is a more challenging task that deserves further investigations in the future.
@@
    Generalization is a universal challenge for state-of-the-art model editing techniques. Except for four instances where FT-L and ROME edit LLMs4Code on the CSNE dataset, achieving a generalization rate surpassing 50%, the generalization rate in all other scenarios dips below 50%.
@@ IV-C RQ3: Improvements of Editing Techniques on LLMs4Code
Specifically, we propose A-GRACE, which introduces an MLP encoder composed of two linear layers. Different from GRACE, the encoder utilizes the mean of all token representations as the key $q\in\mathbb{R}^{d_{n}}$, aiming at capturing the information of the entire sequence from a holistic perspective. The encoding process of the MLP encoder for the key can be described as $q^{*}=W_{2}\sigma(W_{1}q)$, where $W_{1}\in\mathbb{R}^{d_{n}\times\frac{d_{n}}{4}}$, $W_{2}\in\mathbb{R}^{\frac{d_{n}}{4}\times d_{o}}$, $d_{o}$ is the output dimension, and $\sigma$ is the activation function. We set $\sigma$ as ReLu. After that, we train the MLP encoder using contrastive learning on the training set. The training target is to differentiate between GRI and SRI, ensuring that the distance between the keys of the OI and the GRI is as close as possible, while the distance between the keys of the OI and the SRI is as far as possible. To that end, we use the hidden states of the $<$OI, GRI$>$ pairs as positive samples and the hidden states of the $<$OI, SRI$>$ pairs as negative samples for training, with the following loss function [62]:
@@
w/o mean 80.00 91.67 95.11 0.02 9.59 27.03 81.73 97.67 97.81 484.66 87.92 95.68 98.83 0.92 30.84 46.03 83.67 98.03 98.44 510.31
We set the output dimension of the MLP encoder to $d_{o}=256$ to balance encoding performance and efficiency, and the margin to 1.3 in CNLE and 2.5 in CSNE to enable AGRACE to efficiently distinguish between GRI and SRI. During the training phase, we use a batch size of 64, a learning rate of 1e-4, and a weight decay of 1e-4 for 100 epochs. We apply an early stopping strategy with a patience of 10 and use the checkpoint with the best performance on the test set as the final MLP encoder weights.
Results. We set the other parameters of A-GRACE to be consistent with those of GRACE. The evaluation results on the CLMEEval benchmark are shown in Table VI. The results show that A-GRACE consistently outperforms GRACE overall, particularly in terms of the generalization. For instance, when editing CodeLlama on the CNLE dataset, the EM in terms of the generalization is increased from 0.02% of the vanilla GRACE to 64.54% of A-GRACE. On the CSNE dataset, this value is boosted from 9.16% to 97.05%.
@@
Additionally, in this table, “w/o CL” refers to the ablation result of directly using the mean of all token representations as the key without using contrastive learning, and “w/o mean” refers to the ablation result of using the MLP encoder trained by contrastive learning to encode the last token representation. The ablation results indicate that neither directly using the mean of all token representations nor encoding the last token representation with the MLP encoder trained by contrastive learning can improve generalization. In contrast, AGRACE achieves an order of magnitude improvement in generalization without significant declines in other aspects. Notably, on the CSNE dataset, its performance in other aspects is almost identical to that of GRACE. This demonstrates the rationale of our approach design, that is, using a contrastive-learning-augmented MLP encoder to encode the mean of all token representations as the key is the optimal choice.
@@
(1) Using the last token representation as the key cannot help GRACE effectively distinguish between GRI and SRI; (2) By introducing an contrastively-trained encoder in GRACE and encoding the mean of all token representations (the proposed A-GRACE), we successfully enhance the generalization of GRACE without significantly affecting other aspects.